OpenStudy (anonymous):

what is the slope-intercept form of y+4=-2(x-1)

OpenStudy (nincompoop):

solve for Y

OpenStudy (anonymous):

Well first distribute y+4=-2(x-1)  So -2 * x -2 * -1 y + 4 = -2x + 2 Now subtract 4 on both sides to get y by itself y + 4 - 4 = -2x + 2 - 4 = y = -2x -2